OpenStudy (anonymous):

Let r(t) = (24 sin(16t), -4, 24 cos(16t)). Compute d/dt(1/2r'(t) * r"(t))

OpenStudy (anonymous):

r'(t)=(16(24)cos(16t),0,-16(24)sin(16t)) r''(t)=(-16^2(24)sin(16t),0,-16^2(24)cos(16t)) Let w=16^2(24) to keep our sanity r''(t)=(-wsin(16t),0,-wcos(16t)) now before i start the d/dt(1/2)(r'(t)*(r''(t)) is the r'(t) in the denominator or the numerator?

OpenStudy (anonymous):

It's actually 1/2 * r'(t).

OpenStudy (anonymous):

ok then let z=16(24) then r'(t) dot r''(t)= ((z)cos(16t),0,-(z)sin(16t)) * (-(w)sin(16t),0,-(w)cos(16t))= ((-zw)cos(16t)sin(16t)+(zw)cos(16t)sin(16t)) d/dt((r'(t)*r''(t))= Im going to pull out the zw because i already am hating how its looking zw(d/dt(-cos(16t)sin(16t)+sin(16t)cos(16t)) to me thats looking like 0

OpenStudy (anonymous):

which means they are orthogonal (⊥) to eachother

OpenStudy (anonymous):

Even with the 1/2 next to r'(t), will it still equal to 0?

OpenStudy (anonymous):

yes because you can pull the 1/2 out and dot them first then add in the 1/2 after. wz/2 can be pulled out. is the answer not 0?

OpenStudy (anonymous):

Actually I got 0 as well when I was working on this. So I guess it should be right. But I'm still checking on it.

OpenStudy (anonymous):

yeah well its an UGLY problem in deed. well if you took away all coefficents and were left with (sin(t),-4,cos(t))=r(t) r'(t)=(cos(t),0,(-sin(t)) which is ⊥ to r(t) r'"(t)=(-sin(t),0,-cos(t)) which is ⊥ to r'(t)

OpenStudy (anonymous):

r'''(t)=(-cos(t),0,sin(t)) which is ⊥ to r''(t) pattern

OpenStudy (anonymous):

lol, it was annoying and I pretty much got the same as you have. I've asked a couple of my friends about this and they said they all ended up with 0 as well. So maybe that is right.

OpenStudy (anonymous):

Anyways, thanks again for helping me. :)

OpenStudy (anonymous):

welcome